OpenStudy (anonymous):

A line of the form ax+by+c=0 has a gradient of -(2/3) and y intercept at (0,2). Find the integer values of a,b,c. Please solve and explain, thanks :)

OpenStudy (anonymous):

write your equation in term of x. y=(-a/b)x-c/b here gradient = -a/b but also gradient = -2/3 thus a= 2 and b= 3

OpenStudy (anonymous):

given gradient = -2/3 slope of eqn ax+by + c = 0 is -a/b so u can find a,b substitute a and b in the eqn and put x,y values 0,2 to find c

OpenStudy (anonymous):

in the above equation y- intercept = -c/b also y-intercecpt = 2 -c/b=2 put b= 3 and get c